40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Nov 12 01:09:56 2016, gautam, Update, LSC, Recovering DRMI locking 
    Reply  Wed Nov 16 03:10:01 2016, gautam, Update, LSC, DRMI locked on 1f and 3f signals 
       Reply  Wed Nov 16 08:14:43 2016, Steve, Update, LSC, DRMI locked on 1f and 3f signals 5hrs.png
       Reply  Mon Nov 21 14:02:32 2016, gautam, Update, LSC, DRMI locked on 3f signals, arms held on ALS DRMIArms_Nov20.pdf
    Reply  Wed Nov 23 16:21:02 2016, gautam, Update, LSC, ITMY UL glitches are back ITMYproblemsBack.png
       Reply  Mon Nov 28 10:27:13 2016, gautam, Update, SUS, ITMY UL glitches are back ITMY_satboxSpectra.pdfITMY_UL_testerBox.png
       Reply  Tue Nov 29 11:07:37 2016, Steve, Update, LSC, ITMY UL glitches are back gliching400d.png
          Reply  Wed Nov 30 01:47:56 2016, gautam, Update, LSC, Suspension woes IMCwoes.png
             Reply  Wed Nov 30 17:08:56 2016, gautam, Update, LSC, Binary output breakout box removed 
                Reply  Thu Dec 1 02:19:13 2016, gautam, Update, LSC, Binary output breakout box restored MC1_glitch_fast.pngITMY_testerGlitch.pngITMYsatbox.pngULcomparison.pdf
                   Reply  Thu Dec 1 08:02:57 2016, Steve, Update, LSC, glitching ITMY_UL_LL ITMY_UL_LL.png
                      Reply  Fri Dec 2 11:56:42 2016, gautam, Update, LSC, MC1 LEMO jiggled IMG_3474.JPGMC1_LR_dead.png
                         Reply  Mon Dec 5 15:05:37 2016, gautam, Update, LSC, MC1 glitches are back Mc1glitches.png
                      Reply  Thu Dec 8 10:13:43 2016, Steve, Update, LSC, glitching ITMY_UL has a history glitching__ITMY-UL_2007.png
Message ID: 12648     Entry time: Wed Nov 30 01:47:56 2016     In reply to: 12644     Reply to this: 12652
Author: gautam 
Type: Update 
Category: LSC 
Subject: Suspension woes 

Short summary:

  • Looks like Satellite boxes are not to blame for glitchy behaviour of shadow sensor PD readouts
  • Problem may lie at the PD whitening boards (D000210) or with the Contec binary output cards in c1sus
  • Today evening, similar glitchy behaviour was observed in all MC1 PD readout channels, leading to frequent IMC unlocking. Cause unknown, although I did work at 1X5, 1X6 today, and pulled out the PD whitening board for ITMY which sits in the same eurocrate as that for MC1. MC2/MC3 do not show any glitches.

Detailed story below...


Part 1: Satellite box swap

Yesterday, I switched the ITMY and ETMY satellite boxes, to see if the problems we have been seeing with ITMY UL move with the box to ETMY. It did not, while ITMY UL remained glitchy (based on data from approximately 10pm PDT on 28Nov - 10am PDT 29 Nov). Along with the tabletop diagnosis I did with the tester box, I concluded that the satellite box is not to blame.


Part 2: Tracing the signal chain (actually this was part 3 chronologically but this is how it should have been done...)

So if the problem isn't with the OSEMs themselves or the satellite box, what is wrong? I attempted to trace the signal chain from the satellite box into our CDS system as best as I could. The suspension wiring diagram on our wiki page is (I think) a past incarnation. Of course putting together a new diagram was a monumental task I wasn't prepared to undertake tonight, but in the long run this may be helpful. I will put up a diagram of the part I did trace out tomorrow, but the relevant links for this discussion are as follows (? indicates I am unsure):

  1. Sat box (?)--> D010069 via 64pin IDE connector --> D000210 via DB15 --> D990147 via 4pin LEMO connectors --> D080281 via DB25 --> ADC0 of c1sus
  2. D000210 backplane --> cross-connect (mis)labelled "ITMX white" via IDE connector
  3. c1sus CONTEC DO-32L-PE --> D080478 via DB37 --> BO0-1 --> cross-connect labelled "XY220 1Y4-33-16A" via IDE --> (?)  cross-connect (mis)labelled "ITMX white" via IDE connector

I have linked to the DCC page for the various parts where available. Unfortunately I can't locate (on new DCC or old or elog or wiki) drawings for D010069 (Satellite Amplifier Adapter Board), D080281 ("anti-aliasing interface)" or D080478 (which is the binary output breakout box). I have emailed Ben Abbott who may have access to some other archive - the diagrams would be useful as it is looking likely that the problem may lie with the binary output.

So presumably the first piece of electronics after the Satellite box is the PD whitening board. After placing tags on the 3 LEMOs and 1 DB15 cable plugged into this board, I pulled out the ITMY board to do some tabletop diagnosis in the afternoon around 2pm 29Nov.


Part 3: PD whitening board debugging

This particular board has been reported as problematic in the recent past. I started by inserting a tester board into the slot occupied by this board - the LEDs on the tester board suggested that power-supply from the backplane connectors were alright, confirmed with a DMM.

Looking at the board itself, C4 and C6 are tantalum capacitors, and I have faced problems with this type of capacitor in the past. In fact, on the corresponding MC3 board (which is the only one visible, I didn't want to pull out boards unnecessarily) have been replaced with electrolytic capacitors, which are presumably more reliable. In any case, these capacitors do not seem to be at any fault, the board receives +/-15 V as advertised.

The whitening switching is handled by the MAX333 - this is what I looked at next. This IC is essentially a quad SPDT switch, and a binary input supplied via the backplane connector serves to route the PD input either through a whitening filter, or bypass it via a unity gain buffer. The logic levels that effect the switching are +15V and 0V (and not the conventional 5V and 0V), but according to the MAX333 datasheet, this is fine. I looked at the supply voltage to all ICs on the board, DC levels seemed fine (as measured with a DMM) and I also looked at it on an oscilloscope, no glitches were seen in ~30sec viewing stretch. I did notice something peculiar in that with no input supplied to the MAX333 IC (i.e. the logic level should be 15V), the NO and NC terminals appear shorted when checked with a DMM. Zach has noticed something similar in the past, but Koji pointed out that the DMM can be fooled into thinking there is a short. Anyway, the real test was to pull the logic input of the MAX333 to 0, and look at the output, this is what I did next.

The schematic says the whitening filter has poles at 30,100Hz and a zero at 3 Hz. So I supplied as "PD input" a 12Hz 1Vpp sinewave - there should be a gain of ~x4 when this signal passes through the path with the whitening filter. I then applied a low frequency (0.1Hz) square wave (0-5V) to the "bypass" input, and looked at the output, and indeed saw the signal amplitude change by ~4x when the input to the switch was pulled low. This behaviour was confirmed on all five channels, there was no problem. I took transfer functions for all 5 channels (both at the "monitor" point on the backplane connector and on the front panel LEMOs), and they came out as expected (plot to be uploaded soon).

Next, I took the board back to the eurocrate. I first put in a tester box into the slot and measured the voltage levels on the backplane pins that are meant to trigger bypassing of the whitening stage, all the pins were at 0V. I am not sure if this is what is expected, I will have to look inside D080478 as there is no drawing for it. Note that these levels are set using a Contec binary output card. Then I attached the PD whitening board to the tester board, and measured the voltages at the "Input" pins of all the 5 SPDT switches used under 2 conditions - with the appropriate bit sent out via the Contec card set to 0 or 1 (using the button on the suspension MEDM screens). I confirmed using the BIO medm screen that the bit is indeed changing on the software side, but until I look at D080478, I am not sure how to verify the right voltage is being sent out, except to check at the pins on the MAX333. For this test, the UL channel was indeed anomalous - while the other 4 channels yielded 0V (whitening ON, bit=1) and 15V (whitening OFF, bit=0), the corresponding values for the UL channel were 12V and 10V.

I didn't really get any further than this tonight. But this still leaves unanswered questions - if the measured values are faithful, then the UL channel always bypasses the whitening stage. Can this explain the glitchy behaviour?


Part 4: MC1 troublesfrown

At approximately 8pm, the IMC started losing lock far too often - see the attached StripTool trace. There was a good ~2hour stretch before that when I realigned the IMC, and it held lock, but something changed abruptly around 8pm. Looking at the IMC mirror OSEM PD signals, all 5 MC1 channels are glitching frequently. Indeed, almost every IMC lockloss in the attached StripTool is because of the MC1 PD readouts glitching, and subsequently, the damping loops applying a macroscopic drive to the optic which the FSS can't keep up with. Why has this surfaced now? The IMC satellite boxes were not touched anytime recently as far as I am aware. The MC1 PD whitening board sits in the same eurocrate I pulled the ITMY board out of, but squishing cables/pushing board in did not do anything to alleviate the situation. Moreover, MC2 and MC3 look fine, even though their PD whitening boards also sit in the same eurocrate. Because I was out of ideas, I (soft) restarted c1sus and all the models (the thinking being if something was wrong with the Contec boards, a restart may fix it), but there was no improvement. The last longish lock stretch was with the MC1 watchdog turned off, but as soon as I turned it back on the IMC lost lock shortly after.

I am leaving the autolocker off for the night, hopefully there is an easy fix for all of this...

Attachment 1: IMCwoes.png  28 kB  | Hide | Hide all
IMCwoes.png
ELOG V3.1.3-